OpenStudy (anonymous):

After 5 hours of of house painting, Jess asked the assistance of Jon. Together, they finished the job in 4 more hours. Had Jon joined in an hour after Jess started working, the job would have been finished in 7 hours. How long would it take Jess to paint the house alone?

OpenStudy (anonymous):

13 hours

OpenStudy (anonymous):

how?

OpenStudy (dumbcow):

Let J be the rate that jess works, Let T be the rate that they work together 5J +4T = J +6T Solve for T: T = 2J, therefore Jess and Jon work at same rate --> 5J + 4(2J) = 13J House is painted in 13 hours at Jess's rate